# Advances in Robust Federated Learning: Heterogeneity Considerations

### v1

@@ II-A3 Feature Skew
Data feature skew is a kind of data heterogeneity in FL, which refers to the inconsistent distribution of data features on different clients [32, 33], e.g.,$P_{i}(x)\neq P_{j}(x)$. In this scenario, even though the labels are consistent $P_{i}(y)=P_{j}(y)$, the distributions remain different $P_{i}(x|y)\neq P_{j}(x|y)$. As in Fig. 1c, clients $i$, $j$, and $k$ all have the label of dog, but they are photos, cartoons, and art paintings of dogs, respectively. This data feature skew is especially common in Vertical Federated Learning (VFL) scenarios [34, 35]. VFL involves multiple clients that have the same set of sample IDs, but each client holds a different set of features. In other words, each client only knows a portion of the features of the samples. For example, consider two financial institutions, A and B, that collaborate to build a credit scoring model. Institution A holds the basic information and transaction history of its clients, while Institution B holds the loan history and repayment status of its clients. Although these two institutions share the same customer base, they hold different sets of data features. In this longitudinal FL scenario, data feature skew manifest itself in the form of different feature spaces held by agencies A and B.
@@ II-A4 Quality Skew
Data quality skew refers to the difference in data quality on different clients [36, 37]. This can be caused by a variety of factors as shown in Fig. 1d, such as noise in the data acquisition process, the data processing flow of different clients, or the accuracy of data labeling. Data quality skew can have a significant impact on the performance of FL models [38]. In this scenario, some clients may have high-quality data that has been accurately measured and carefully labeled, while others may contain a lot of noise or inaccurate labels. High-quality data can provide more accurate information that helps the model learn effective feature representations and prediction rules. On the contrary, low-quality data may introduce misleading information that causes the model to learn incorrect feature relationships, thus reducing the model’s generalization ability and prediction accuracy [39].
@@ II-A5 Quantity Skew
Data quantity skew refers to a significant difference in the amount of data on different clients [40, 41]. For example, some clients may have a large number of data samples while others have only a small amount of data as shown in Fig. 1e. This quantity skew affects the training process of a FL model because the model may tend to optimize the performance on clients with a larger amount of data while ignoring clients with a smaller amount of data. The impact of data quantity skew on the model is mainly in two aspects [42]: on the one hand, clients with larger data quantities may have a larger impact on the global model update, which results in the model performing better on these clients and worse on clients with smaller data quantities. On the other hand, the imbalance in the amount of data may result in the model not being able to fully learn the features of the data on all clients, thus reducing the generalization ability of the model.
@@ II-B1 Model Structure Heterogeneity
Model structure heterogeneity refers to the fact that different clients use different model structures to adapt to their respective data characteristics and computational resources [44]. This heterogeneity mainly stems from the diversity of devices and the diversity of task requirements in the real world. For example, in a FL system, some clients may be servers equipped with high-performance GPUs capable of handling complex deep learning models such as deep convolutional neural networks (CNNs) or transformer models for image recognition or natural language processing tasks. While other clients may be resource-constrained edge devices, such as smartphones or IoT sensors, which may only be able to run lightweight models, such as shallow neural networks or decision trees, to reduce computational and storage overhead.
Moreover, even for the same type of task, different clients may choose different model structures based on the characteristics of their local data [45]. For example, in the task of performing handwritten digit recognition, some clients may use a standard multilayer perceptron (MLP) model, while others may employ a convolutional neural network model with spatial feature extraction capabilities. This structural difference results in different dimensions and properties of the model parameters, making it difficult to directly apply traditional FL algorithms, such as the Federated Averaging Algorithm (FedAvg), which typically assume that all clients use the same model structure.
@@ II-C Task Heterogeneity
In FL, task heterogeneity refers to the fact that clients involved in the learning process may face different task requirements or objective functions [9]. This heterogeneity may stem from the environment in which the clients are located, differences in user behavior, or specific application scenario requirements, resulting in different objective functions or learning tasks that need to be optimized for each client.
For example [48], in a FL application in the medical domain, different hospitals may focus on different types of disease diagnosis, so their learning tasks and objective functions may be different, e.g., Fig. 3a demonstrates that clients have homogeneous models but the distributions have labels for different medicines. Fig. 3b demonstrates when clients have similar datasets such as face dataset, and individual clients have heterogeneous classification labels. Fig. 3c Task types are heterogeneous, including classification and regression tasks.
Task heterogeneity introduces additional complexity to FL because traditional FL frameworks typically assume that all devices jointly optimize a global objective function [49]. To cope with task heterogeneity, FL systems need to employ more flexible algorithms and policies that allow different devices to personalize model training based on their own task requirements. In addition, task heterogeneity also requires that FL systems are able to effectively coordinate the learning processes of different devices to ensure a balance between the generalization capabilities of the global model and the individual needs of each client. Such coordination may require appropriate adaptation of the model updating and parameter aggregation processes to suit the characteristics and needs of different tasks.
Figure 3: Illustration of federated task heterogeneity.
@@ II-D Communication heterogeneity
In FL, communication heterogeneity refers to the differences in communication capabilities of the devices involved in the learning process [10, 50]. Such differences are mainly in terms of network bandwidth, latency and connection stability, which directly affect the transmission efficiency of model parameters and the synchronization of the whole learning process. For example, devices are usually deployed in different network environments with different network connection settings (3G, 4G, 5G, Wi-Fi) [51, 52]. Devices with high bandwidth and low latency can quickly upload and download model parameters, thus completing a round of model updates in a short period of time. On the contrary, devices with limited bandwidth and high latency may face longer waiting time when transmitting model parameters, resulting in the training progress of the whole system being affected. The heterogeneity of communication increases the communication cost and complexity to some extent [53]. In addition, differences in connection stability can also have an impact on the efficiency of FL. Unstable connections may result in lost or delayed model parameter updates, which in turn affects the convergence speed and accuracy of the model.
@@ II-E Device heterogeneity
In the application scenario of FL, device heterogeneity is a crucial issue that cannot be neglected [54, 55]. Due to the wide variety of client devices participating in FL, there are significant differences in computing power, storage resources, energy supply, etc., which leads to device heterogeneity [11]. This heterogeneity leads to inconsistent efficiency and capability of different devices in processing data and training models, which may affect the performance and efficiency of the whole FL system. For example, high-performance devices such as servers or high-end smartphones are able to process complex models and large data volumes quickly due to their powerful processors and ample storage space, thus taking on more computational burden during FL and speeding up the entire learning process [56]. On the contrary, low-performance devices, such as low-end smartphones, are less capable of model training and data processing due to the limitations of computational resources and storage space, and may also slow down the progress of the entire FL system, which may become the bottleneck of system efficiency. This is because the system usually needs to wait for all devices to complete their local updates before global model aggregation can take place. In addition, devices with weaker computational power may face greater pressure on energy consumption when processing complex models, which is especially critical in energy-constrained environments.
Therefore, FL algorithms and system design need to take into account the heterogeneity of computational capabilities among devices, and ensure efficient operation and fairness of the system through strategies such as reasonable task allocation, model simplification, and resource optimization to ensure that all devices can effectively participate in the learning process, while maintaining the coordination and fairness of the overall system.
@@ III-A1 Data Cleaning
Li et al. [57] presents an optimization problem i.e., selecting a high-quality set of training samples for a FL task under a given budget in a privacy-preserving manner without knowledge of the participants’ local data and training process. Ma et al. [58] proposes a federated data cleaning protocol, called FedClean, for edge intelligence scenarios that aims to achieve data cleaning without compromising data privacy. FedClean Different edge nodes first generate a Boolean share of their data and distribute it to two clean servers. These two servers then run the FedClean protocol to compute and sort the data entry scores. Faced with the problem of noisy data labels, Chen et al. [59] proposes a robust FL by considering two networks with different data distributions. The core idea is that the edge devices learn to assign local weights of the loss function in the noisy labeled dataset and cooperate with the central server to update the global weights as a way to help edge nodes that do not have a clean dataset to reweight their noisy loss functions. Similarly, Duan et al. et al. [60] proposes a new data filtering approach to deal with label noise in joint learning, called FED-DR-Filter, which focuses on identifying clean data by exploiting the correlation of global data representations. Tsouvalas et al. [61] proposes FedLN framework enough to deal with labeling noise in different FL training phases, i.e., FL initialization, on-device model training, and server model aggregation, which is able to adapt to the different computational power of the devices in the FL system. Li et al. [62] proposed FedDiv to address the challenge of FL with noisy labels, and they designed a global noise filter with a joint noise filter for efficiently identifying samples with noisy labels on each client, which improves the stability of the local training process. Xu et al. [63] proposed a collaborative data filtering method Safe to deal with poisoning attacks, using the alternating direction method of multipliers to detect the attacked devices whose training process will be interrupted, and using the K-mean clustering algorithm to cluster the trusted data to find out and filter out the poisoned data from the attacked devices.
@@ III-A3 Representation Learning
To address the problem of data heterogeneity, Liu et al. [71] introduces federated transfer learning (FTL). FTL does not compromise user privacy and can transfer knowledge across domains in a data federation. This enables the target domain to utilize the rich labels of the source domain to build flexible and efficient models. Li et al. [72] proposes FedMask, a FL framework for efficient communication and computation, where each device can learn an efficient personalized structured sparse model. Zhang et al. [73] proposes for Federated Unsupervised Representation Learning (FURL) to learn generic representation models without supervision while preserving data privacy. FURL solves the problem of inconsistent and misaligned representation space through a dictionary module and an alignment module. Similarly, Yan et al. [74] uses unsupervised representation to pre-train deep neural networks in a federated environment using unlabeled data, which allows for the extraction of discriminative features using pre-trained networks. Wu et al. [75] proposes a communication-efficient FL algorithm, FedOnce, which is based on the privacy-preserving technique of momentary accounting and employs an unsupervised learning representation that enables better results to be achieved with only one-time communication between the parties. Li et al. [76] proposed MOON to correct for local training of individuals using similarities between model representations. Jang et al. [77] proposes a personalized federated learning approach that applies local feature representation learning to stabilize the decision boundary and improve the local feature extraction capability of the client. Tan et al. [78] proposes FedProto, which only passes the prototype to the client without passing model parameters and gradients.
@@ III-B Model-Level Methods
In this section, we discuss some advanced model-level FL methods for solving various heterogeneity types. Model-level methods means that these methods improve FL in terms of model parameters, model output, etc. Specifically, we classify these methods based on their adopted techniques, including parameter decoupling, model compression, knowledge distillation and federated optimization.
@@ III-B1 Parameter Decoupling
Parameter Decoupling based methods regard the whole model parameters as the composition of components, which play different roles in the process of training or inferring. Depending on the goal of parameter decoupling, different methods usually have different granularities. Some methods only regard the whole model as the composition of a backbone model and a head model, while some others decouple the model in a layer-wise or channel-wise, even coordinate-wise manner. Arivazhagan et al. [79] propose FedPer, which split global model into base layers and personalized layers. Base layers are layers closer to the input side while the personalized layers are the remaining layers, and each client can customize the own architecture of personalized layers while sharing the architecture of base layers. In each round, only the parameters of base layers are uploaded to the server and then aggregated. Based on the belief that data from all parties involved in FL often share global feature representations while reflect distribution and label skew in the labels, Collins et al. [80] propose FedRep, which separate model into representation part and head part. The clients train a global representation in a federated manner and then train their own head based on the learned representation. Luo et al. [81] empirically verify that distribution and label skew impact more on deeper layers, and then propose CCVR. CCVR utilizes regularization and calibration in local training to reduce weight bias, and trains a GMM to generate virtual representations for further classifier training. To solve the problem of long-tailed data in FL, Shang et al. [82] propose CReFF, where clients upload real feature gradients, and then the server use them to optimize federated feature gradients for re-training classifier. Comparing to above works that enable shared global representations of features, Liang et al. [83] propose LG-FedAvg, where clients share a common architecture of classifier, and each client learns compressed local representation for training a classifier suitable for local representations of all clients. Learning such local representations on the clients can also help alleviating task heterogeneity, encouraging the clients to learn and extract representations suitable for global classifiers. Xu et al. [84] propose FedPAC, which encourage each client to learn global task variant information by regularizing local representations to approach the global feature centroid. In order to train personalized client models under feature skew, distribution skew and other scenarios, Shen et al. [85] propose CD2-pFed, which implement channel-wised personalization by channel-decoupling, and utilize personalized-weights to dynamically control the degree of personalization in each layer. Besides, CD2-pFed also utilizes cyclic-distillation to mutually distill local and global model representations, bridging the semantic gap between the decoupled channels. Recently, Chen et al. [86] suggest that biased projection heads may cause unreliability on models when faced with data heterogeneity. Therefore, they propose APH method, where clients can get multiple initialized parameters from existing projection heads, and then fine-tune and average them, thus eliminating the bias.
@@
In order to reduce the memory cost of clients, Pillutla et al. [87] propose to divide the model parameters into two parts: shared parameters and personal parameters. Therefore, they propose FedSim and FedAlt, representing training both part of parameters in a simultaneous manner or an alternate manner, respectively. Yi et al. [45] propose personalized FL method FedGH, where clients share a generalized head. The clients upload local average representation, which is utilized for individually training of the generalized head on the server side. Recently, to achieve model personalization when it is model heterogeneity, Wang et al. [88] propose pFedHR. In each round, the server performs layer-wise decomposition on client’s models, and then groups the layers. After that, the server uses the grouped layers to reassembly candidate models and fine-tune them with a public dataset.
@@
To alleviate the impact of device heterogeneity and communication heterogeneity, Diao et al. [51] propose HeteroFL, where the server determines the complexity level of each client based on its computation and communication capabilities, and assigns sub-models of different sizes to clients according to their complexity level. Recently, Yi et al. [89] point out that even pruned models can also reveal the real architecture of the global model. Therefore, they propose a novel framework FedP3, where clients only upload selected model parameters back to the server for privacy protection. Pfeiffer et al. [90] empirically observe that the techniques based on sub-models suffer from performance degradation because the co-adaptation between parameters is broken. Thus, they propose Successive Layer Training (SLT). All clients can train the same parameters, but they begin at early layers with a down-scale head. After early layers are well trained, then SLT squeeze the parameters of early layers and expand the head layers for further training.

### v2

@@ II-A3 Feature Skew
Data feature skew is the inconsistent distribution of data features across clients [84], e.g., $P_{i}(x)\neq P_{j}(x)$. In this scenario, even though the labels are consistent $P_{i}(y)=P_{j}(y)$, the distributions remain different $P_{i}(x|y)\neq P_{j}(x|y)$. As in Fig. 1c, clients $i$, $j$, and $k$ all have the label of dog, but they are photos, cartoons, and art paintings of dogs, respectively. This data feature skew is especially common in Vertical FL (VFL) scenarios [85]. VFL involves multiple clients that have the same set of sample IDs, but each client holds a different set of features. In other words, each client only knows a portion of the features of the samples.
@@ II-A4 Quality Skew
Data quality skew refers to the difference in data quality on different clients [86]. This can be caused by a variety of factors as shown in Fig. 1d, such as noise in the data collection process or errors in the labelling of the collected data. In this scenario, some clients may have high-quality data that has been accurately measured and carefully labeled, while others may contain a lot of noise or inaccurate labels. High-quality data can provide more accurate information that helps the model learn effective feature representations and prediction rules. On the contrary, low-quality data may introduce misleading information that causes the model to learn incorrect feature relationships, thus reducing the model’s generalization ability and prediction accuracy [45].
@@ II-A5 Quantity Skew
Data quantity skew refers to a significant difference in the amount of data on different clients [87], as shown in Fig. 1e. This quantity skew affects the training process of a FL model because the model may tend to optimize the performance on clients with a larger amount of data while ignoring clients with a smaller amount of data. The impact of data quantity skew on the model is mainly in two aspects [88]: on the one hand, clients with larger data quantities may have a larger impact on the global model update, which results in the model performing better on these clients and worse on clients with smaller data quantities. On the other hand, the imbalance in the amount of data may result in the model not being able to fully learn the features of the data on all clients, thus reducing the generalization ability of the model.
@@ II-B1 Model Structure Heterogeneity
Model structure heterogeneity refers to the fact that different clients use different model structures to adapt to their respective data characteristics and computational resources [89]. This heterogeneity mainly stems from the diversity of devices and the diversity of task requirements in the real world. For example, in a FL system, some clients may be servers equipped with high-performance GPUs capable of handling complex deep learning models such as CNNs or transformer models for image recognition or natural language processing tasks. While other clients may be resource-constrained edge devices, such as smartphones or IoT sensors, which may only be able to run lightweight models, such as shallow neural networks or decision trees, to reduce computational and storage overhead
Moreover, even for the same type of task, different clients may choose different model structures based on the characteristics of their local data [50]. For example, in the task of performing handwritten digit recognition, some clients may use a standard multilayer perceptron (MLP) model, while others may employ a convolutional neural network model with spatial feature extraction capabilities. This structural difference results in different dimensions and properties of the model parameters, making it difficult to directly apply traditional FL algorithms, such as FedAvg, which typically assume that all clients use the same model structure.
@@ II-C Task Heterogeneity
In FL, task heterogeneity refers to the fact that clients involved in the learning process may face different task requirements or objective functions [10]. This heterogeneity may stem from the environment in which the clients are located, differences in user behavior, or specific application scenario requirements, resulting in different objective functions or learning tasks that need to be optimized for each client. For example [92], in a FL application in the medical domain, different hospitals may focus on different types of disease diagnosis, so their learning tasks and objective functions may be different, e.g., Fig. 2a demonstrates that clients have homogeneous models but the distributions have labels for different medicines. Fig. 2b demonstrates when clients have similar datasets such as face dataset, and individual clients have heterogeneous classification labels. Fig. 2c Task types are heterogeneous, including classification and regression tasks.
Task heterogeneity introduces additional complexity to FL because traditional FL frameworks typically assume that all devices jointly optimize a global objective function [93]. Task heterogeneity requires FL systems to be able to efficiently coordinate the learning process across devices to ensure a balance between the generalisation capabilities of the global model and the personalised needs of the client. Such coordination may require appropriate adaptation of the model updating and parameter aggregation processes to suit the characteristics and needs of different tasks [94].
Figure 2: FL task heterogeneity.
@@ II-D Communication heterogeneity
Communication heterogeneity refers to the differences in communication capabilities of the devices involved in the learning process [11]. Such differences are mainly in terms of network bandwidth, latency and connection stability, which directly affect the transmission efficiency of model parameters and the synchronization of the whole learning process. For example, devices are usually deployed in different network environments with different network connection settings (3G, 4G, 5G, Wi-Fi) [56, 95], as shown in Fig. 5. Devices with high bandwidth and low latency can quickly upload and download model parameters, thus completing a round of model updates in a short period of time. On the contrary, devices with limited bandwidth and high latency may face longer waiting time when transmitting model parameters, resulting in the training progress of the whole system being affected. The heterogeneity of communication increases the communication cost and complexity to some extent [96]. In addition, differences in connection stability can also have an impact on the efficiency of FL. Unstable connections may result in lost or delayed model parameter updates, which in turn affects the convergence speed and accuracy of the model. For example, in Fig. 5, though the middle client accesses the network through 5G connection, the connection is unstable and has a higher risk of packet loss, which can hinder the process of FL.
Figure 3: FL model heterogeneity.
Figure 4: FL communication heterogeneity.
Figure 5: FL device heterogeneity.
@@ II-E Device heterogeneity
In the application scenario of FL, device heterogeneity is a crucial issue that cannot be neglected [97]. Due to the wide variety of client devices participating in FL, there are significant differences in computing power, storage resources, energy supply, etc., which leads to device heterogeneity [12]. As shown in Fig. 5, each client is equipped with a different type of device, resulting in a significant performance span between the clients. This heterogeneity leads to inconsistent efficiency and capability of different devices in processing data and training models, which may affect the performance and efficiency of the whole FL system. For example, high-performance devices such as servers or high-end smartphones are able to process complex models and large data volumes quickly due to their powerful processors and ample storage space, thus taking on more computational burden during FL and speeding up the entire learning process [98]. On the contrary, low-performance devices, such as low-end smartphones, are less capable of model training and data processing due to the limitations of computational resources and storage space, and may also slow down the progress of the entire FL system, which may become the bottleneck of system efficiency. This is because the system usually needs to wait for all devices to complete their local updates before global model aggregation can take place. In addition, devices with weaker computational power may face greater pressure on energy consumption when processing complex models, which is especially critical in energy-constrained environments.
@@ III-A1 Data Cleaning
Li et al. [99] presents an optimization problem i.e., selecting a high-quality set of training samples for a FL task under a given budget in a privacy-preserving manner without knowledge of the participants’ local data and training process. Faced with the problem of noisy data labels, Chen et al. [100] proposes a robust FL by considering two networks with different data distributions. The core idea is that the edge devices learn to assign local weights of the loss function in the noisy labeled dataset and cooperate with the central server to update the global weights as a way to help edge nodes that do not have a clean dataset to reweight their noisy loss functions. Similarly, Duan et al. et al. [101] proposes a new data filtering approach to deal with label noise in joint learning, called FED-DR-Filter, which focuses on identifying clean data by exploiting the correlation of global data representations. Tsouvalas et al. [102] proposes FedLN framework enough to deal with labeling noise in different FL training phases, i.e., FL initialization, on-device model training, and server model aggregation, which is able to adapt to the different computational power of the devices in the FL system. Li et al. [44] proposed FedDiv to address the challenge of FL with noisy labels, and they designed a global noise filter with a joint noise filter for efficiently identifying samples with noisy labels on each client, which improves the stability of the local training process. Xu et al. [103] proposed a collaborative data filtering method Safe to deal with poisoning attacks, using the alternating direction method of multipliers to detect the attacked devices whose training process will be interrupted, and using the K-mean clustering algorithm to cluster the trusted data to find out and filter out the poisoned data from the attacked devices.
@@ III-A3 Representation Learning
To address the problem of data heterogeneity, Liu et al. [39] introduces federated transfer learning, which does not compromise user privacy and can transfer knowledge across domains in a data federation. This enables the target domain to utilize the rich labels of the source domain to build flexible and efficient models. Li et al. [55] proposes FedMask, a FL framework for efficient communication and computation, where each device can learn an efficient personalized structured sparse model. Zhang et al. [110] proposes for Federated Unsupervised Representation Learning (FURL) to learn generic representation models without supervision while preserving data privacy. FURL solves the problem of inconsistent and misaligned representation space through a dictionary module and an alignment module. Similarly, Yan et al. [111] uses unsupervised representation to pre-train deep neural networks in a federated environment using unlabeled data, which allows for the extraction of discriminative features using pre-trained networks. Wu et al. [40] proposes a communication-efficient FL algorithm, FedOnce, which is based on the privacy-preserving technique of momentary accounting and employs an unsupervised learning representation that enables better results to be achieved with only one-time communication between the parties. Li et al. [25] proposed MOON to correct for local training of individuals using similarities between model representations, and thus reduce the impact of data heterogeneity. Jang et al. [112] proposes a personalized FL approach that applies local feature representation learning to stabilize the decision boundary and improve the local feature extraction capability of the client. Tan et al. [49] proposes FedProto, which only passes the prototype to the client without passing model parameters and gradients, which can be available when it is data heterogeneity and model heterogeneity.
@@ III-B Model-Level Methods
Model-level methods means improving FL by adopting some techniques that only apply to model parameters, model outputs, etc., and make minor changes to the input data and FL architecture. Specifically, we classify these methods based on their techniques, including parameter decoupling, model compression, knowledge distillation and federated optimization.
@@ III-B1 Parameter Decoupling
Parameter Decoupling based methods regard the whole model parameters as the composition of components, which play different roles in the process of training or inferring. Depending on the goal of parameter decoupling, different methods usually have different granularities. Some methods only regard the whole model as the composition of a backbone model and a head model, while some others decouple the model in a layer-wise or channel-wise, even coordinate-wise manner. Arivazhagan et al. [54] propose FedPer, which split global model into base layers and personalized layers. Base layers are layers closer to the input side while the personalized layers are the remaining layers. Personalized layers is customized while base layers are shared and transmitted. Therefore, FedPer can also to some extent handles task heterogeneity and communication heterogeneity. Based on the belief that data from all parties involved in FL often share global feature representations while reflect distribution and label skew in the labels, Collins et al. [37] propose FedRep, which separate model into representation part and head part. The clients train a global representation in a federated manner and then train their own head based on the learned representation. Luo et al. [26] empirically verify that distribution and label skew impact more on deeper layers, CCVR, which utilizes regularization and calibration in local training to reduce weight bias, and trains a GMM to generate virtual representations for further classifier training, which can help reducing the impact of data heterogeneity. To solve the problem of long-tailed data in FL, Liang et al. [47] propose LG-FedAvg, where clients share a common architecture of classifier, and each client learns compressed local representation for training a classifier suitable for local representations of all clients. Learning such local representations on the clients can also help alleviating task heterogeneity, encouraging the clients to learn and extract representations suitable for global classifiers. Xu et al. [28] propose FedPAC, which encourage each client to learn global task variant information by regularizing local representations to approach the global feature centroid. Besides, it also trains personalized classifiers and collaborates similar clients, reducing inter-client feature variance. In order to train personalized client models under feature skew, distribution skew and other scenarios, Shen et al. [42] propose CD2-pFed, which implement channel-wised personalization by channel-decoupling, and utilize personalized-weights to dynamically control the degree of personalization in each layer. Besides, CD2-pFed also utilizes cyclic-distillation to mutually distill local and global model representations, bridging the semantic gap between the decoupled channels.
@@
Some works based on parameters decoupling are also proposed for handling model heterogeneity. To reduce the memory cost of clients, Pillutla et al. [113] propose to divide the model parameters into two parts: shared parts and personal parts. Therefore, they propose FedSim and FedAlt, representing training both part of parameters in a simultaneous manner or an alternate manner, respectively. Yi et al. [50] propose model heterogeneous method FedGH, where clients share a generalized head. The clients upload local average representation, which is utilized for individually training of the generalized head on the server side. However, the above works require some specific parts of the model structure to be identical among the clients, which can’t be used for the scenarios with thorough model heterogeneity. Recently, to achieve personalization when models are heterogeneous, Wang et al. [52] propose pFedHR. The server decomposes and then groups the client’s models layer by layer in each round. Afterwards, the server reassembles the candidate models using the grouped layers and fine-tunes them using a public dataset.
@@
Parameter decoupling can also be utilized for handling device heterogeneity, which is performed by reducing the size of the model. To alleviate the impact of device heterogeneity and communication heterogeneity, Diao et al. [56] propose HeteroFL, where the server determines the complexity level of each client based on its computation and communication capabilities, and assigns sub-models of different sizes to clients according to their complexity level. However, comparing to traditional FL methods, HeteroFL suffers from significant performance degradation. Recently, Yi et al. [62] point out that even pruned models can also reveal the real architecture of the global model. Therefore, they propose a novel framework FedP3, where clients only upload selected model parameters back to the server for privacy protection. Pfeiffer et al. [57] empirically observe that the techniques based on sub-models suffer from performance degradation because the co-adaptation between parameters is broken. Thus, they propose SLT. All clients can train the same parameters, but they begin at early layers with a down-scale head. After early layers are well trained, then SLT squeeze the parameters of early layers and expand the head layers for further training.